OpenStudy (anonymous):

Simplify t^2 - 4t - 21 / t + 3

hero (hero):

t^2-7t+3t -21/t+3) = t(t-7)+3(t-7)/(t+3) = (t-7)(t+3)/(t+3) = (t-7)

OpenStudy (turingtest):

\[(t^2-4t-21)=(t+3)(t-7)\]\[[(t+3)(t-7)]/(t+3)=t-7\]